Assume the Cobb-Douglas production function
is Q= L ^0.75 K^0.5 and if price of labor per day is 5 birr and price of capital per day is 10 birr; and if total outlay (cost budget) per day is 400 birr,
A.Find L and K that maximize out put
B.What is the maximum out put the equilibrium L* and K*
1
Expert's answer
2021-08-16T08:47:21-0400

a)

Given 

"Q =L^{0.75} K^{0.5} ......... (1)"

Price of labor = 5 birr per day 

Price of capital = 10 birr per day 

Total cost budget = 400 birr per day

Budget line or isocost line 

"5L+10K=400 ......(2)"


Maximize "Q =L^{0.75} K^{0.5}"

Subject to constraints "5L+10K=400"

Using lagrangian method

"X =L^{0.75} K^{0.5} +\u03bb(400-5L-10K)"

FOC

"\frac{dX}{dL}=0, \frac{dX}{dK}=0, \frac{dX}{dλ}=0"

Differentiating Lagrangian function with respect to L

"\frac{dX}{dL}=0\\⇒0.75L^{−0.25}K^{0.5}=5λ ..........(3)\\
Now\\

\frac{dX}{dK}=0\\⇒0.5L^{0.75}K^{−0.5}=10λ ..........(4)\\
and \\

\frac{dX}{dλ}=0\\⇒5L+10K=400 ..........(2)"

Dividing eq 3 and 4

"⇒\frac{0.75L^{−0.25}K^{0.5}}{0.5L^{0.75}K^{−0.5}}=\frac{5λ}{10λ}\\⇒\frac{3 K}{2L}=\frac{1}{2}\\⇒K=\frac{L}{3}"

Putting value of K in eq 2

"5L+10(\frac{L}{3}) = 400\\⇒\frac{15+10}{3}L=400\\⇒L=400×\frac{3}{25}\\⇒L=48\\ 
And\space K = \frac{48}{3} = 16"

The optimal level of L is 48 and K is 16 that maximizes the output. 


(b)

Maximum output will be

"Q^*=48^{0.75}×16^{0.5}\\⇒Q^*= 72.944"